A Complete Bibliography of Publications in Technometrics for the decade 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

Analyses [You59]. Analysis [Ans59, Bir59, Man59a, Pla59]. Analytical [Dav59]. Application [Ano59g, Bud59]. Aspects [Dav59]. Averages [Rob59].

Back [Ano59a, Ano59b, Ano59c, Ano59d]. Balance [Ano59g, Ans59, Bud59, Sat59]. Based [Rob59]. Budne [YKT^{+}59].

Calculations [BH59]. Censored [Coh59]. Certain [Man59a]. Chart [Rob59]. Chemical [You59]. Column [Man59a]. compact [Tuk59]. Condensed [BH59]. Control [Jac59, Rob59].

Data [Pla59]. Designs [Ano59g, Ano59f, Bud59, De 59]. Discussion [YKT^{+}59]. Distribution [Coh59]. Duckworth [Tuk59]. Duplication [Dyk59].

Economics [Dav59]. Editor [Ano59e]. Electron [Kao59]. Errata [Ano59g, Ano59f, Ano60]. Estimation [Kao59]. Estimators [Coh59]. Evaluating [Kah59]. Evaluation [You59]. Evolutionary [BH59]. Experimentation [Sat59]. Experiments [Ano60, Ans59, Bir59, Dan59, Dyk59, Zel59].

Factorial [Ano60, Bir59, Dan59, Dyk59, Zel59]. Factors [Ano59f, De 59]. Front [Ano59h, Ano59i, Ano59j, Ano59k].

Geometric [Rob59]. Graphical [Kao59].

Half [Dan59]. Half-Normal [Dan59]. Hazardous [Kah59].

Information [Ano59l]. Interaction [Man59a]. Interpreting [Dan59].

Known [YCS59].

LaGrange [US59]. Latin [Man59a]. Level [Dan59]. Levels [Ano59f, De 59]. Life [Ano60, Kao59, Pla59, Zel59]. Life-Testing [Kao59].

Made [YCS59]. Manufacturing [Kah59]. Matching [YCS59]. Mathematical [Fis59]. Matter [Ano59a, Ano59b, Ano59c, Ano59d, Ano59h, Ano59i, Ano59j, Ano59k]. Measurements [YCS59]. Measuring [Man59b]. Messrs. [YKT^{+}59]. Methods [Ans59, Jac59]. Mixed [Kao59]. Model [Kah59]. Moving [Rob59]. Multipliers [US59].

Natural [Fis59]. Normal [Coh59, Dan59]. Note [Ano59e].

Operation [BH59].

papers [YKT^{+}59]. Parameters [Kao59]. Partial [Dyk59]. Plants [Kah59]. Plots [Dan59]. Probability [Fis59]. Process [Man59b]. Products [Kah59]. Programs [BH59].

Quality [Jac59]. Quick [Ans59, Tuk59].

Random [Ano59g, Ans59, Bud59, Sat59]. Related [Jac59]. Reliability [Kah59]. Replication [Bir59]. Response [Ano59f, De 59, US59]. Rocks [You59]. Row [Man59a]. Row-Column [Man59a].

Safety [Kah59]. sample [Tuk59]. Samples [Coh59]. Sciences [Fis59]. Screening [Ans59]. Several [Jac59]. Simplified [Coh59]. Singly [Coh59]. Some [Dav59]. specifications [Tuk59]. Squares [Man59a]. Standards [YCS59]. Statistical [Dav59, Kah59]. Surface [Ano59f, De 59]. Surfaces [US59]. Systems [Kah59].

Test [Pla59, Tuk59]. Testing [Ano60, Dav59, Kao59, Zel59]. Tests [Rob59]. Three [Ano59f, De 59]. Truncated [Coh59]. Tubes [Kao59]. Two [Dan59, You59, Tuk59]. Two-Level [Dan59]. two-sample [Tuk59]. Type [Man59a].

Use [Dan59, US59].

Variables [Jac59]. Volume [Ano59l].

Weibull [Kao59]. without [Bir59].

References

Anonymous:1959:BMa

[Ano59a] Anonymous. Back matter. Technometrics, 1(1):97–100, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

         tronic). URL http://www.jstor.org/stable/1266312.

Anonymous:1959:BMb

[Ano59b] Anonymous. Back matter. Technometrics, 1(2):211–216, May 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

         URL http://www.jstor.org/stable/1266470.

Anonymous:1959:BMc

[Ano59c] Anonymous. Back matter. Technometrics, 1(3):309–310, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

         tronic). URL http://www.jstor.org/stable/1266448.

Anonymous:1959:BMd

[Ano59d] Anonymous. Back matter. Technometrics, 1(4):420–424, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

         (electronic). URL http://www.jstor.org/stable/1266723.

Anonymous:1959:EN

[Ano59e] Anonymous. Editor’s note. Technometrics, 1(2):??, May 1959. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

         URL http://www.jstor.org/stable/1266464.

Anonymous:1959:ERS

[Ano59f] Anonymous. Errata: “Response Surface Designs for Three Factors at Three Levels”. Technometrics, 1(4):419, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266722. See [De 59].

Anonymous:1959:EAR

[Ano59g] Anonymous. Errata: “The Application of Random Balance Designs”. Technometrics, 1(4):419, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266721. See [Bud59].

Anonymous:1959:FMa

[Ano59h] Anonymous. Front matter. Technometrics, 1(1):??, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266304.

Anonymous:1959:FMb

[Ano59i] Anonymous. Front matter. Technometrics, 1(2):??, May 1959. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266463.

Anonymous:1959:FMc

[Ano59j] Anonymous. Front matter. Technometrics, 1(3):??, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266441.

Anonymous:1959:FMd

[Ano59k] Anonymous. Front matter. Technometrics, 1(4):??, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266714.

Anonymous:1959:VI

[Ano59l] Anonymous. Volume information. Technometrics, 1(4):425–426, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266713.

Anonymous:1960:EFE

[Ano60] Anonymous. Errata: “Factorial Experiments in Life Testing”. Technometrics, 2(1):121, February 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266537. See [Zel59].

Anscombe:1959:QAM

[Ans59] F. J. Anscombe. Quick analysis methods for random balance screening experiments. Technometrics, 1(2):195–209, May 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266469.

Box:1959:CCE

[BH59] G. E. P. Box and J. S. Hunter. Condensed calculations for evolutionary operation programs. Technometrics, 1(1):77–95, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266311.

Birnbaum:1959:AFE

[Bir59] Allan Birnbaum. On the analysis of factorial experiments without replication. Technometrics, 1(4):343–357, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266716.

Budne:1959:ARB

[Bud59] Thomas A. Budne. The application of random balance designs. Technometrics, 1(2):139–155, May 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266467. See discussion [YKT^{+}59] and errata [Ano59g].

Cohen:1959:SEN

[Coh59] A. Clifford Cohen, Jr. Simplified estimators for the normal distribution when samples are singly censored or truncated. Technometrics, 1(3):217–237, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266442.

Daniel:1959:UHN

[Dan59] Cuthbert Daniel. Use of half-normal plots in interpreting factorial two-level experiments. Technometrics, 1(4):311–341, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266715.

Davies:1959:SSA

[Dav59] O. L. Davies. Some statistical aspects of the economics of analytical testing. Technometrics, 1(1):49–61, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266309.

DeBaun:1959:RSD

[De 59] Robert M. De Baun. Response surface designs for three factors at three levels. Technometrics, 1(1):1–8, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266305. See errata [Ano59f].

Dykstra:1959:PDF

[Dyk59] O. Dykstra, Jr. Partial duplication of factorial experiments. Technometrics, 1(1):63–75, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266310.

Fisher:1959:MPN

[Fis59] Ronald A. Fisher. Mathematical probability in the natural sciences. Technometrics, 1(1):21–29, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266307.

Jackson:1959:QCM

[Jac59] J. Edward Jackson. Quality control methods for several related variables. Technometrics, 1(4):359–377, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266717.

Kahn:1959:SME

[Kah59] Louis B. Kahn. A statistical model for evaluating the reliability of safety systems for plants manufacturing hazardous products. Technometrics, 1(3):293–307, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266447.

Kao:1959:GEM

[Kao59] John H. K. Kao. A graphical estimation of mixed Weibull parameters in life-testing of electron tubes. Technometrics, 1(4):389–407, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1266719.

Mandel:1959:ALS

[Man59a] John Mandel. The analysis of Latin squares with a certain type of row-column interaction. Technometrics, 1(4):379–387, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1266718.

Mandel:1959:MP

[Man59b] John Mandel. The measuring process. Technometrics, 1(3):251–267, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266444.

Plackett:1959:ALT

[Pla59] R. L. Plackett. The analysis of life test data. Technometrics, 1(1):9–19, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1266306.

Roberts:1959:CCT

[Rob59] S. W. Roberts. Control chart tests based on geometric moving averages. Technometrics, 1(3):239–250, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266443.

Satterthwaite:1959:RBE

[Sat59] F. E. Satterthwaite. Random balance experimentation. Technometrics, 1(2):111–137, May 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266466. See discussion [YKT^{+}59].

Tukey:1959:QCT

[Tuk59] John W. Tukey. A quick, compact, two-sample test to Duckworth’s specifications. Technometrics, 1(1):31–48, February 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1266308.

Umland:1959:ULM

[US59] A. W. Umland and W. N. Smith. The use of LaGrange multipliers with response surfaces. Technometrics, 1(3):289–292, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1266446.

Youden:1959:MMM

[YCS59] W. J. Youden, W. S. Connor, and N. C. Severo. Measurements made by matching with known standards. Technometrics, 1(2):101–109, May 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1266465.

Youden:1959:DPM

[YKT^{+}59] W. J. Youden, Oscar Kempthorne, John W. Tukey, G. E. P. Box, J. S. Hunter, F. E. Satterthwaite, and Thomas A. Budne. Discussion of the papers of Messrs. Satterthwaite and Budne. Technometrics, 1(2):157–193, May 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266468. See [Sat59, Bud59].

Youden:1959:ECA

[You59] W. J. Youden. Evaluation of chemical analyses on two rocks. Technometrics, 1(4):409–417, November 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1266720.

Zelen:1959:FEL

[Zel59] Marvin Zelen. Factorial experiments in life testing. Technometrics, 1(3):269–288, August 1959. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266445. See errata [Ano60].